A far sighted person has his near point 50cm,find the power of lens he should use to see at 25cm, clearlly
1
Expert's answer
2019-06-11T10:01:14-0400

according to the formula of the lens:

"D=\frac{1}{d}-\frac{1}{f} = \frac{1}{0.25}-\frac{1}{0.5} = 2 dioptriy"